1ReviewI Constants Periodic Table Part A A NaT ion moves from inside a cell, where the electric potential is -70 mV, to outsid23ReviewI Constants I Periodic Table Each of the following vectors is given in terms of its x- and y components. z 23 m/s, ty 3
Answer #1

Only one question can be posted as per HOMEWORKLIB RULES

1.

1 mV=10-3 V

Change in Ion's Potential energy

2.

a)

Answers are

  • the distance between the positive charge and the electron
  • the charge of the positive charge
  • The charge on the electron

Since electric field is given by

E=KQ/r2

-----------------------------------------------------------------------------------------------

To find acceleration ,electric field is required to find acceleration .Here full question not provided

The magnitude of acceleration is

a=|q|E/m

-------------------------------------------------------------------------------

3.

a)

V=sqrt[232+352] =41.88 m/s

b)

o=tan-1(35/23) =56.69o=56.7o

c)

a=sqrt[4.52+(-6.5)2] =7.91 m/s2

d)

o=tan-1(-6.5/4.5) =-55.3o

o=55.3o below +x-axis